# Characteristic Learning for Provable One Step Generation

### v3

@@ C.1. Auxiliary Properties
(E.5) $\|\widehat{Z}(t_{k})-\widehat{Z}_{k}\|_{2}\leq\frac{(1+\|{\mathscr{B}}\|_{\operatorname{Lip}}\tau)^{k}-1}{\|{\mathscr{B}}\|_{\operatorname{Lip}}}D\kappa(T)R\tau,\quad 1\leq k\leq K.$
We prove (E.5) by induction. When $k=1$, since that $\widehat{Z}(0)=\widehat{Z}_{0}$, it follows from (E.4) that
@@
which satisfies (E.5). For general $k\geq 2$ we assume that (E.5) holds up to $k-1$. Then applying (E.4) implies that
@@
which proves that (E.5) is true for $k$. Therefore, we have verified the inequality (E.5). Substituting $K\tau=T$ and $(1+\|{\mathscr{B}}\|_{\operatorname{Lip}}\tau)^{k}\leq\exp(\|{\mathscr{B}}\|_{\operatorname{Lip}}k\tau)$into (E.5) completes the proof. ∎
@@
Combining Lemmas D.3 and E.1, we have
@@
Taking expectation on both sides of the inequality with respect to $X_{0}$ and $X_{1}$ and recalling the definition of 2-Wasserstein distance (1.2) implies (E.6).
@@
where we used Lemma E.6. Combining this with Theorem 3.11 implies the desired result. ∎
Appendix F Proof of Results in Section 3.5
In this section, we aim to prove Theorem 3.13. Towards this end, we first relate the averaged $2$-Wasserstein distance of the characteristic generator to its $L^{2}$-risk by Lemma F.1. Then an oracle inequality of $L^{2}$-risk are proposed in Lemma F.2. Finally, by substituting approximation and generalization error bounds into the oracle inequality and using Theorem 3.11 completes the proof.
Lemma F.1.
@@
Lemma F.2 (Oracle inequality for characteristic fitting).
@@
${\mathbb{E}}_{{\mathcal{Z}}}\big{[}{\mathcal{R}}_{T}(\widehat{g})\big{]}$ $\leq\inf_{g\in{\mathscr{G}}}{\mathcal{R}}_{T,R}(g)+CW_{2}^{2}(\widehat{\mu}_{K},\mu_{T})$
@@
Next we introduce a spatial truncation to (F.2), which implies the following risk:
@@
We then define the following semi-discretized risk, which replaces the empirical average with respect to first two variables in (F.3) by its population :
(F.4) ^$\widehat{{\mathcal{R}}}_{T,R,m}(g)=\frac{1}{m}\sum_{i=1}^{m}\Big{\{}\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}\|g^{*}(t,s,Z_{t}^{(i)})-g(t,s,Z_{t}^{(i)})\|_{2}^{2}\mathbbm{1}\{\|Z_{t}^{(i)}\|_{\infty}\leq R\}\,\mathrm{d}s\mathrm{d}t\Big{\}}.$
@@
(F.5) ${\mathcal{R}}_{T}(g)={\mathbb{E}}_{Z_{0}\sim\mu_{0}}\Big{[}\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}\|g^{*}(t,s,Z_{t})-g(t,s,Z_{t})\|_{2}^{2}\,\mathrm{d}s\mathrm{d}t\Big{]},$
@@
(F.6) ${\mathcal{R}}_{T,R}(g)={\mathbb{E}}_{Z_{0}\sim\mu_{0}}\Big{[}\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}\|g^{*}(t,s,Z_{t})-g(t,s,Z_{t})\|_{2}^{2}\mathbbm{1}\{\|Z_{t}\|_{\infty}\leq R\}\,\mathrm{d}s\mathrm{d}t\Big{]}.$
According to definitions (F.1) to (F.6), it is straightforward that for each $g\in{\mathscr{G}}$,
${\mathcal{R}}_{T}(\widehat{g})$ $\leq\Big{(}{\mathcal{R}}_{T}(\widehat{g})-{\mathcal{R}}_{T,R}(\widehat{g})\Big{)}+\Big{(}{\mathcal{R}}_{T,R}(\widehat{g})-2\widehat{{\mathcal{R}}}_{T,R,m}(\widehat{g})\Big{)}+2\Big{(}\widehat{{\mathcal{R}}}_{T,R,m}(\widehat{g})-\widehat{{\mathcal{R}}}_{T,R,m,K}(\widehat{g})\Big{)}$
@@
where the inequality follows from the fact that $\widehat{g}$ is the minimizer of the empirical risk minimizer ^$\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}$ over the hypothesis class ${\mathscr{G}}$. Taking expectation on both sides of the inequality with respect to ${\mathcal{Z}}\sim\mu_{0}^{m}$ yields
(F.7) ${\mathbb{E}}_{{\mathcal{Z}}}\big{[}{\mathcal{R}}_{T}(\widehat{g})\big{]}$ $\leq\Big{(}{\mathcal{R}}_{T}(\widehat{g})-{\mathcal{R}}_{T,R}(\widehat{g})\Big{)}+{\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\sup_{g\in{\mathscr{G}}}{\mathcal{R}}_{T,R}(g)-2\widehat{{\mathcal{R}}}_{T,R,m}(g)\Big{]}$
$\quad+2{\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\widehat{{\mathcal{R}}}_{T,R,m}(\widehat{g})-\widehat{{\mathcal{R}}}_{T,R,m,K}(\widehat{g})\Big{]}+2{\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\widehat{{\mathcal{R}}}_{T,R,m,K}(\widehat{g})-\widehat{{\mathcal{R}}}_{T,m,K}(\widehat{g})\Big{]}$
@@
Up to now, the $L^{2}$-risk of the estimator $\widehat{g}$ is divided into six terms by (F.7). In the remainder of the proof, we bound six these error terms one by one.
@@
    The first term in the right-hand side of (F.7) measures the error caused by the spatial truncation, for which
    (F.8) ${\mathcal{R}}_{T}(\widehat{g})-{\mathcal{R}}_{T,R}(\widehat{g})\leq\frac{CR^{2}}{\exp(\theta R^{2})},$
@@
    The second term in the right-hand side of (F.7) is known as the generalization error, for which
    (F.9) ${\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\sup_{g\in{\mathscr{G}}}{\mathcal{R}}_{T,R}(g)-2\widehat{{\mathcal{R}}}_{T,R,m}(g)\Big{]}\leq CR^{2}\max_{1\leq k\leq d}\frac{\operatorname{VCdim}(\Pi_{k}{\mathscr{G}})}{m\log^{-1}(m)},$
@@
    The third term in the right-hand side of (F.7) is led by the time discretization. It holds for each $g\in{\mathscr{G}}$ that
    (F.10) ^$\widehat{{\mathcal{R}}}_{T,R,m}(g)-\widehat{{\mathcal{R}}}_{T,R,m,K}(g)\leq\frac{CR^{2}}{K},$
@@
    The fourth term in the right-hand side of (F.7) is also a truncation error. By an argument similar to (F.8), it holds that
    (F.11) ${\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\widehat{{\mathcal{R}}}_{T,R,m,K}(\widehat{g})-\widehat{{\mathcal{R}}}_{T,m,K}(\widehat{g})\Big{]}\leq\frac{CR^{2}}{\exp(\theta R^{2})},$
@@
    The fifth term in the right-hand side of (F.7) is caused by the error of Euler method, for which
    (F.12) ${\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\widehat{{\mathcal{R}}}_{T,m,K}(\widehat{g})-2\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}(\widehat{g})\Big{]}\leq CW_{2}^{2}(\widehat{\mu}_{K},\mu_{T}),$
@@
    The sixth term in the right-hand side of (F.7) is the empirical risk of the estimator. Using the definition of the empirical risk minimizer, we deduce
    (F.13) ${\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}(g)\Big{]}$ $\leq\inf_{g\in{\mathscr{G}}}{\mathcal{R}}_{T,R}(g)+CW_{2}^{2}(\widehat{\mu}_{K},\mu_{T})+\frac{CR^{2}}{K}+\frac{CR^{2}}{\exp(\theta R^{2})},$
@@
Plugging (F.8) to (F.13) into (F.7) obtains the desired result.
Step 1. Estimate the first term in the right-hand side of (F.7).
For each hypothesis $g\in{\mathscr{G}}$, by an argument similar to (D.5), we have
@@
(F.14) $\leq 8\Big{(}{\mathbb{E}}_{Z_{t}\sim\mu_{t}}^{1/2}\Big{[}\|Z_{s}\|_{2}^{4}\Big{]}+{\mathbb{E}}_{Z_{t}\sim\mu_{t}}^{1/2}\Big{[}\|g(t,s,Z_{t})\|_{2}^{4}\Big{]}\Big{)}\operatorname{Pr}^{1/2}\{\|Z_{t}\|_{\infty}>R\},$
@@
$Z_{s}\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}X_{s}\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\alpha_{s}X_{0}+\beta_{s}U+\sigma\beta_{s}\epsilon,\quad U\sim\nu,\leavevmode\nobreak\ \epsilon\sim N(0,I_{d}),$
which implies by an argument similar to (D.7) that
@@
where the last inequality follows from Lemma A.7. As a consequence,
(F.15) ${\mathbb{E}}_{Z_{t}}^{1/2}\Big{[}\|g^{*}(t,s,Z_{t})\|_{2}^{4}\Big{]}={\mathbb{E}}_{Z_{s}}^{1/2}\Big{[}\|Z_{s}\|_{2}^{4}\Big{]}\leq C,$
@@
(F.16) ${\mathbb{E}}_{Z_{t}}^{1/2}\Big{[}\|g(t,s,Z_{t})\|_{2}^{4}\Big{]}\leq dB_{\operatorname{flow}}^{2}R^{2}.$
Further, using (D.8) yields
(F.17) $\sup_{t\in(0,1)}\operatorname{Pr}\big{\{}\|Z_{t}\|_{\infty}>R\big{\}}=\sup_{t\in(0,1)}\operatorname{Pr}\big{\{}\|X_{t}\|_{\infty}>R\big{\}}\leq\frac{2d}{\exp(\theta R^{2})},$
where $\theta$ is a constant only depending on $\sigma$. Substituting inequalities (F.15), (F.16) and (F.17) into (F.14) deduces
@@
where $C$ is a constant only depending to $d$ and $\sigma$. Finally, combining the above inequality with definitions (F.5) and (F.6) completes the proof of (F.8).
Step 2. Estimate the second term in the right-hand side of (F.7).
For simplicity of notation, we define the $k$-th term of ${\mathcal{R}}_{T,R}$ (F.6) and ^$\widehat{{\mathcal{R}}}_{T,R,m}$ (F.4), respectively, as
@@
^$\widehat{{\mathcal{R}}}_{T,R,m}^{k}(b)$ $=\frac{1}{m}\sum_{i=1}^{m}\Big{\{}\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}(g_{k}^{*}(t,s,Z_{t}^{(i)})-g_{k}(t,s,Z_{t}^{(i)}))^{2}\mathbbm{1}\{\|Z_{t}^{(i)}\|_{\infty}\leq R\}\,\mathrm{d}s\mathrm{d}t\Big{\}}.$
Applying the boundedness of $g\in{\mathscr{G}}$, Proposition 3.3, and Lemma G.1 yields
@@
Summing the above inequalities with respect to $1\leq k\leq d$ completes the proof of (F.9).
Step 3. Estimate the third term in the right-hand side of (F.7).
@@
$u(t,s,x)=\|g^{*}(t,s,x)-\widehat{g}(t,s,x)\|_{2}^{2},\quad 0\leq t\leq s\leq T,\leavevmode\nobreak\ x\in{\mathbb{B}}_{R}^{\infty}.$
@@
(F.18) $|\partial_{t}u(t,s,x)|\leq 2\|g^{*}(t,s,x)-g(t,s,x)\|_{2}\|\partial_{t}g^{*}(t,s,x)-\partial_{t}g(t,s,x)\|_{2}\leq CR^{2},$
@@
(F.19) $|\partial_{s}u(t,s,x)|\leq CR^{2},\quad 0\leq t\leq s\leq T,\leavevmode\nobreak\ x\in{\mathbb{B}}_{R}^{\infty}.$
Substituting (F.18) and (F.19) into Lemma F.3 yields (F.10).
Step 4. Estimate the forth term in the right-hand side of (F.7).
We use an argument similar to Step 1. For each $10\leq k\leq\ell\leq K-$ 1, it follows that
@@
(F.20) $=8\Big{(}{\mathbb{E}}_{{\mathcal{Z}}}^{1/2}\Big{[}\|Z_{\ell\tau}^{(i)}\|_{2}^{4}\Big{]}+{\mathbb{E}}_{{\mathcal{Z}}}^{1/2}\Big{[}\|\widehat{g}(k\tau,\ell\tau,Z_{k\tau}^{(i)})\|_{2}^{4}\Big{]}\Big{)}\operatorname{Pr}^{1/2}\big{\{}\|Z_{k\tau}^{(i)}\|_{\infty}>R\big{\}},$
where we used Cauchy-Schwarz inequality and the triangular inequality. Substituting inequalities (F.15), (F.16) and (F.17) into (F.20) yields that for $10\leq k\leq\ell\leq K-$ 1,
(F.21) ${\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\|g^{*}(k\tau,\ell\tau,Z_{k\tau}^{(i)})-\widehat{g}(k\tau,\ell\tau,Z_{k\tau}^{(i)})\|_{2}^{2}\mathbbm{1}\{\|Z_{k\tau}^{(i)}\|_{\infty}>R\}\Big{]}\leq\frac{CR^{2}}{\exp(\theta R^{2})},$
where $C$ is a constant only depending on $d$ and $\sigma$. Combining (F.21) with definitions (F.2) and (F.3) deduces (F.11).
Step 5. Estimate the fifth term in the right-hand side of (F.7).
For each $1\leq i\leq m$ and $10\leq k\leq\ell\leq K-$ 1, it follows that
@@
(F.22) $\leq 4\|\widehat{Z}_{\ell}^{(i)}-Z_{\ell\tau}^{(i)}\|_{2}^{2}+4\|\widehat{g}\|_{\operatorname{Lip}}^{2}\|\widehat{Z}_{k}^{(i)}-Z_{k\tau}^{(i)}\|_{2}^{2}+2\|g^{*}(k\tau,\ell\tau,Z_{k\tau}^{(i)})-\widehat{g}(k\tau,\ell\tau,Z_{k\tau}^{(i)})\|_{2}^{2}.$
Substituting (F.22) into (F.1) deduces
^$\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}(\widehat{g})$ $\leq\frac{2}{mK^{2}}\sum_{i=1}^{m}\sum_{k=0}^{K-1}\Big{\{}\frac{1}{2}\Big{(}4\|\widehat{Z}_{k}^{(i)}-Z_{k\tau}^{(i)}\|_{2}^{2}+4\|\widehat{g}\|_{\operatorname{Lip}}^{2}\|\widehat{Z}_{k}^{(i)}-Z_{k\tau}^{(i)}\|_{2}^{2}\Big{)}$
$\quad+\sum_{\ell=k+1}^{K-1}\Big{(}4\|\widehat{Z}_{\ell}^{(i)}-Z_{\ell\tau}^{(i)}\|_{2}^{2}+4\|\widehat{g}\|_{\operatorname{Lip}}^{2}\|\widehat{Z}_{k}^{(i)}-Z_{k\tau}^{(i)}\|_{2}^{2}\Big{)}\Big{\}}+2\widehat{{\mathcal{R}}}_{T,m,K}(\widehat{g})$
$\leq 4(1+\|\widehat{g}\|_{\operatorname{Lip}}^{2})\frac{1}{m}\sum_{i=1}^{m}\|\widehat{Z}_{K}^{(i)}-Z_{T}^{(i)}\|_{2}^{2}+2\widehat{{\mathcal{R}}}_{T,m,K}(\widehat{g}),$
@@
^$\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}(\widehat{g})-2\widehat{{\mathcal{R}}}_{T,m,K}(\widehat{g})\leq 4(1+\|\widehat{g}\|_{\operatorname{Lip}}^{2})\frac{1}{m}\sum_{i=1}^{m}\|\widehat{Z}_{K}^{(i)}-Z_{T}^{(i)}\|_{2}^{2}.$
@@
(F.23) ${\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}(\widehat{g})-2\widehat{{\mathcal{R}}}_{T,m,K}(\widehat{g})\Big{]}\leq CW_{2}^{2}(\widehat{\mu}_{K},\mu_{T}),$
where $C$ is a constant only depending on $d$ and $\sigma$. By the same argument as inequality (F.23), we can obtain (F.12) immediately.
Step 6. Estimate the sixth term in the right-hand side of (F.7).
@@
where the first term can be estimated by (F.23), the second and third terms can be bounded by an argument similar to (F.11) and (F.10), respectively. For the last term, we have ${\mathbb{E}}_{{\mathcal{Z}}}[\widehat{{\mathcal{R}}}_{T,R,m}(g)]={\mathcal{R}}_{T,R}(g)$. Combining above results yields (F.13). ∎

### v5

@@ D.1. Auxiliary Properties
(F.5) $\|\widehat{Z}(t_{k})-\widehat{Z}_{k}\|_{2}\leq\frac{(1+\|{\mathscr{B}}\|_{\operatorname{Lip}}\tau)^{k}-1}{\|{\mathscr{B}}\|_{\operatorname{Lip}}}D\kappa(T)R\tau,\quad 1\leq k\leq K.$
We prove (F.5) by induction. When $k=1$, since that $\widehat{Z}(0)=\widehat{Z}_{0}$, it follows from (F.4) that
@@
which satisfies (F.5). For general $k\geq 2$ we assume that (F.5) holds up to k1k-1. Then applying (F.4) implies that
@@
which proves that (F.5) is true for $k$. Therefore, we have verified the inequality (F.5). Substituting $K\tau=T$ and $(1+\|{\mathscr{B}}\|_{\operatorname{Lip}}\tau)^{k}\leq\exp(\|{\mathscr{B}}\|_{\operatorname{Lip}}k\tau)$into (F.5) completes the proof. ∎
@@
Combining Lemmas E.3 and F.1, we have
@@
Taking expectation on both sides of the inequality with respect to $X_{0}$ and $X_{1}$ and recalling the definition of 2-Wasserstein distance implies (F.6).
@@
where we used Lemma F.6. Combining this with Theorem 3.12 implies the desired result. ∎
Appendix G Proof of Results in Section 3.5
In this section, we aim to prove Theorem 3.14. Towards this end, we first relate the averaged $2$-Wasserstein distance of the characteristic generator to its $L^{2}$-risk by Lemma G.1. Then an oracle inequality of $L^{2}$-risk are proposed in Lemma G.2. Finally, by substituting approximation and generalization error bounds into the oracle inequality and using Theorem 3.12 completes the proof.
Lemma G.1.
@@
Lemma G.2 (Oracle inequality for characteristic fitting).
@@
[…]
@@
Next we introduce a spatial truncation to (G.2), which implies the following risk:
@@
We then define the following semi-discretized risk, which replaces the empirical average with respect to first two variables in (G.3) by its population :
(G.4) $\widehat{{\mathcal{R}}}_{T,R,m}(g)=\frac{1}{m}\sum_{i=1}^{m}\Big\{\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}\|g^{*}(t,s,Z_{t}^{(i)})-g(t,s,Z_{t}^{(i)})\|_{2}^{2}\mathbbm{1}\{\|Z_{t}^{(i)}\|_{\infty}\leq R\}\,\mathrm{d}s\mathrm{d}t\Big\}.$
@@
(G.5) ${\mathcal{R}}_{T}(g)={\mathbb{E}}_{Z_{0}\sim\mu_{0}}\Big[\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}\|g^{*}(t,s,Z_{t})-g(t,s,Z_{t})\|_{2}^{2}\,\mathrm{d}s\mathrm{d}t\Big],$
@@
(G.6) ${\mathcal{R}}_{T,R}(g)={\mathbb{E}}_{Z_{0}\sim\mu_{0}}\Big[\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}\|g^{*}(t,s,Z_{t})-g(t,s,Z_{t})\|_{2}^{2}\mathbbm{1}\{\|Z_{t}\|_{\infty}\leq R\}\,\mathrm{d}s\mathrm{d}t\Big].$
According to definitions (G.1) to (G.6), it is straightforward that for each $g\in{\mathscr{G}}$,
[…]
@@
where the inequality follows from the fact that $\widehat{g}$ is the minimizer of the empirical risk minimizer $\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}$ over the hypothesis class $\mathscr{G}$}. Taking expectation on both sides of the inequality with respect to ${\mathcal{Z}}\sim\mu_{0}^{m}$ yields
[…]
@@
Up to now, the $L^{2}$-risk of the estimator $\widehat{g}$ is divided into six terms by (G.7). In the remainder of the proof, we bound six these error terms one by one.
@@
    The first term in the right-hand side of (G.7) measures the error caused by the spatial truncation, for which
    (G.8) ${\mathcal{R}}_{T}(\widehat{g})-{\mathcal{R}}_{T,R}(\widehat{g})\leq\frac{CR^{2}}{\exp(\theta R^{2})},$
@@
    The second term in the right-hand side of (G.7) is known as the generalization error, for which
    (G.9) ${\mathbb{E}}_{{\mathcal{Z}}}\Big[\sup_{g\in{\mathscr{G}}}{\mathcal{R}}_{T,R}(g)-2\widehat{{\mathcal{R}}}_{T,R,m}(g)\Big]\leq CR^{2}\max_{1\leq k\leq d}\frac{\operatorname{VCdim}(\Pi_{k}{\mathscr{G}})}{m\log^{-1}(m)},$
@@
    The third term in the right-hand side of (G.7) is led by the time discretization. It holds for each $g\in{\mathscr{G}}$ that
    (G.10) $\widehat{{\mathcal{R}}}_{T,R,m}(g)-\widehat{{\mathcal{R}}}_{T,R,m,K}(g)\leq\frac{CR^{2}}{K},$
@@
    The fourth term in the right-hand side of (G.7) is also a truncation error. By an argument similar to (G.8), it holds that
    (G.11) ${\mathbb{E}}_{{\mathcal{Z}}}\Big[\widehat{{\mathcal{R}}}_{T,R,m,K}(\widehat{g})-\widehat{{\mathcal{R}}}_{T,m,K}(\widehat{g})\Big]\leq\frac{CR^{2}}{\exp(\theta R^{2})},$
@@
    The fifth term in the right-hand side of (G.7) is caused by the error of Euler method, for which
    (G.12) ${\mathbb{E}}_{{\mathcal{Z}}}\Big[\widehat{{\mathcal{R}}}_{T,m,K}(\widehat{g})-2\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}(\widehat{g})\Big]\leq CW_{2}^{2}(\widehat{\mu}_{K},\mu_{T}),$
@@
    The sixth term in the right-hand side of (G.7) is the empirical risk of the estimator. Using the definition of the empirical risk minimizer, we deduce
    (G.13) ${\mathbb{E}}_{{\mathcal{Z}}}\Big[\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}(g)\Big]$ $\leq\inf_{g\in{\mathscr{G}}}{\mathcal{R}}_{T,R}(g)+CW_{2}^{2}(\widehat{\mu}_{K},\mu_{T})+\frac{CR^{2}}{K}+\frac{CR^{2}}{\exp(\theta R^{2})},$
@@
Plugging (G.8) to (G.13) into (G.7) obtains the desired result.
Step 1. Estimate the first term in the right-hand side of (G.7).
For each hypothesis $g\in{\mathscr{G}}$, by an argument similar to (E.5), we have
@@
(G.14) $\leq 8\Big({\mathbb{E}}_{Z_{t}\sim\mu_{t}}^{1/2}\Big[\|Z_{s}\|_{2}^{4}\Big]+{\mathbb{E}}_{Z_{t}\sim\mu_{t}}^{1/2}\Big[\|g(t,s,Z_{t})\|_{2}^{4}\Big]\Big)\operatorname{Pr}^{1/2}\{\|Z_{t}\|_{\infty}>R\},$
@@
$Z_{s}\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}X_{s}\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\alpha_{s}X_{0}+\beta_{s}U+\sigma\beta_{s}\epsilon,\quad U\in\nu,\penalty 10000\ \epsilon\sim N(0,I_{d}),$
which implies by an argument similar to (E.7) that
@@
where the last inequality follows from Lemma B.7. As a consequence,
(G.15) ${\mathbb{E}}_{Z_{t}}^{1/2}\Big[\|g^{*}(t,s,Z_{t})\|_{2}^{4}\Big]={\mathbb{E}}_{Z_{s}}^{1/2}\Big[\|Z_{s}\|_{2}^{4}\Big]\leq C,$
@@
(G.16) ${\mathbb{E}}_{Z_{t}}^{1/2}\Big[\|g(t,s,Z_{t})\|_{2}^{4}\Big]\leq dB_{\operatorname{flow}}^{2}R^{2}.$
Further, using (E.8) yields
(G.17) $\sup_{t\in(0,1)}\operatorname{Pr}\big\{\|Z_{t}\|_{\infty}>R\big\}=\sup_{t\in(0,1)}\operatorname{Pr}\big\{\|X_{t}\|_{\infty}>R\big\}\leq\frac{2d}{\exp(\theta R^{2})},$
where $\theta$ is a constant only depending on $\sigma$. Substituting inequalities (G.15), (G.16) and (G.17) into (G.14) deduces
@@
where $C$ is a constant only depending to $d$ and $\sigma$. Finally, combining the above inequality with definitions (G.5) and (G.6) completes the proof of (G.8).
Step 2. Estimate the second term in the right-hand side of (G.7).
For simplicity of notation, we define the $k$-th term of ${\mathcal{R}}_{T,R}$ (G.6) and $\widehat{{\mathcal{R}}}_{T,R,m}$ (G.4), respectively, as
@@
$\widehat{{\mathcal{R}}}_{T,R,m}^{k}(b)$ $=\frac{1}{m}\sum_{i=1}^{m}\Big\{\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}(g_{k}^{*}(t,s,Z_{t}^{(i)})-g_{k}(t,s,Z_{t}^{(i)}))^{2}\mathbbm{1}\{\|Z_{t}^{(i)}\|_{\infty}\leq R\}\,\mathrm{d}s\mathrm{d}t\Big\}.$
Applying the boundedness of $g\in{\mathscr{G}}$, Proposition 3.3, and Lemma I.1 yields
@@
Summing the above inequalities with respect to $1\leq k\leq d$ completes the proof of (G.9).
Step 3. Estimate the third term in the right-hand side of (G.7).
@@
$u(t,s,x)=\|g^{*}(t,s,x)-\widehat{g}(t,s,x)\|_{2}^{2},\quad 0\leq t\leq s\leq T,\penalty 10000\ x\in{\mathbb{B}}_{R}^{\infty}.$
@@
(G.18) $|\partial_{t}u(t,s,x)|\leq 2\|g^{*}(t,s,x)-g(t,s,x)\|_{2}\|\partial_{t}g^{*}(t,s,x)-\partial_{t}g(t,s,x)\|_{2}\leq CR^{2},$
@@
(G.19) $|\partial_{s}u(t,s,x)|\leq CR^{2},\quad 0\leq t\leq s\leq T,\penalty 10000\ x\in{\mathbb{B}}_{R}^{\infty}.$
Substituting (G.18) and (G.19) into Lemma G.3 yields (G.10).
Step 4. Estimate the forth term in the right-hand side of (G.7).
We use an argument similar to Step 1. For each $0\leq k\leq\ell\leq K-1$, it follows that
@@
(G.20) $=8\Big({\mathbb{E}}_{{\mathcal{Z}}}^{1/2}\Big[\|Z_{\ell\tau}^{(i)}\|_{2}^{4}\Big]+{\mathbb{E}}_{{\mathcal{Z}}}^{1/2}\Big[\|\widehat{g}(k\tau,\ell\tau,Z_{k\tau}^{(i)})\|_{2}^{4}\Big]\Big)\operatorname{Pr}^{1/2}\big\{\|Z_{k\tau}^{(i)}\|_{\infty}>R\big\},$
where we used Cauchy-Schwarz inequality and the triangular inequality. Substituting inequalities (G.15), (G.16) and (G.17) into (G.20) yields that for $0\leq k\leq\ell\leq K-1$,
(G.21) ${\mathbb{E}}_{{\mathcal{Z}}}\Big[\|g^{*}(k\tau,\ell\tau,Z_{k\tau}^{(i)})-\widehat{g}(k\tau,\ell\tau,Z_{k\tau}^{(i)})\|_{2}^{2}\mathbbm{1}\{\|Z_{k\tau}^{(i)}\|_{\infty}>R\}\Big]\leq\frac{CR^{2}}{\exp(\theta R^{2})},$
where $C$ is a constant only depending on $d$ and $\sigma$. Combining (G.21) with definitions (G.2) and (G.3) deduces (G.11).
Step 5. Estimate the fifth term in the right-hand side of (G.7).
For each $1\leq i\leq m$ and $0\leq k\leq\ell\leq K-1$, it follows that
@@
(G.22) $\leq 4\|\widehat{Z}_{\ell}^{(i)}-Z_{\ell\tau}^{(i)}\|_{2}^{2}+4\|\widehat{g}\|_{\operatorname{Lip}}^{2}\|\widehat{Z}_{k}^{(i)}-Z_{k\tau}^{(i)}\|_{2}^{2}+2\|g^{*}(k\tau,\ell\tau,Z_{k\tau}^{(i)})-\widehat{g}(k\tau,\ell\tau,Z_{k\tau}^{(i)})\|_{2}^{2}.$
Substituting (G.22) into (G.1) deduces
[…]
@@
$\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}(\widehat{g})-2\widehat{{\mathcal{R}}}_{T,m,K}(\widehat{g})\leq 4(1+\|\widehat{g}\|_{\operatorname{Lip}}^{2})\frac{1}{m}\sum_{i=1}^{m}\|\widehat{Z}_{K}^{(i)}-Z_{T}^{(i)}\|_{2}^{2}.$
@@
(G.23) ${\mathbb{E}}_{{\mathcal{Z}}}\Big[\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}(\widehat{g})-2\widehat{{\mathcal{R}}}_{T,m,K}(\widehat{g})\Big]\leq CW_{2}^{2}(\widehat{\mu}_{K},\mu_{T}),$
where $C$ is a constant only depending on $d$ and $\sigma$. By the same argument as inequality (G.23), we can obtain (G.12) immediately.
Step 6. Estimate the sixth term in the right-hand side of (G.7).
@@
where the first term can be estimated by (G.23), the second and third terms can be bounded by an argument similar to (G.11) and (G.10), respectively. For the last term, we have ${\mathbb{E}}_{{\mathcal{Z}}}[\widehat{{\mathcal{R}}}_{T,R,m}(g)]={\mathcal{R}}_{T,R}(g)$. Combining above results yields (G.13). ∎